# Gaze3P: Gaze-Based Prediction of User-Perceived Privacy

### v1

@@ Abstract.
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In this work, we explore a novel approach: The use of human eye gaze as an implicit source of information on user-perceived privacy. Prior research has shown that eye gaze contains rich information about the user, such as identifiers (Cantoni et al., 2015), quasi-identifiers (Sammaknejad et al., 2017), confidential attributes (e.g. user activities (Steil and Bulling, 2015), attentive (Faber et al., 2018; Vertegaal et al., 2003) and cognitive states (Huang et al., 2016; Bulling and Roggen, 2011), or information about private situations (Steil et al., 2019)). Consequently, we hypothesise that gaze data also provides insights into users’ perceived privacy levels. To this end, we present Gaze3P – the first large-scale dataset for studying user-perceived privacy from the perspective of human eye gaze. Gaze3P includes gaze data (i.e. where, when, and how a person looks) of $100$ participants viewing $1000$ natural images showing different objects, some with private attributes (e.g. credit cards and medical history) as shown in Fig. 2. The dataset also provides user ratings of perceived privacy on a scale from 1 (very private) to 7 (very safe) for each image. The full dataset, including all annotations, will be made publicly available upon acceptance (cf. (Author, s) for the implementation).
@@
Using our new dataset, we then explore different learning tasks, focusing on the automatic prediction of users’ perceived privacy solely from gaze behaviour using machine learning. These tasks include: Stimuli-based Perceived Privacy (SPP) tasks to infer how private a stimulus (e.g. image) is and User-based Perceived Privacy (UPP) tasks to infer informtion about the user (e.g. privacy expertise or identity). Our models demonstrate that human eye gaze provides accurate predictions of perceived privacy.
@@
We use the Gaze3P predictions of user-perceived privacy to optimise the parameters of differentially private mechanisms. Differentially private mechanisms obfuscate sensitive data samples such that only a limited amount of information about the private data can still be deduced from the obfuscated output of the mechanism. The exact amount of acceptable leakage depends on a privacy budget parameter $\varepsilon>0$, which determines the obfuscating noise added by the mechanism. If $\varepsilon$ is small, the privacy guarantee becomes stronger, but usually, the output of the mechanism is less accurate, and usability decreases. It is therefore important to not choose the privacy budget $\varepsilon$ too small, i.e. to only add the minimal amount of noise that guarantees a target privacy level. The optimisation of DP-parameters has therefore seen much attention in recent years (Hsu et al., 2014; Lee and Clifton, 2011; Jorgensen et al., 2015; Boenisch et al., 2022).
@@
Our new dataset Gaze3P and the resulting ML model predictions provide a new way to determine $\varepsilon$ which reflects a user’s perceived privacy. Depending on the actual use case we propose different mappings from perceived privacy levels to $\varepsilon$-values. We evaluate how each mapping affects the utility of the obfuscated output dataset and show that our gaze-based approach outperforms previous work.
@@
    We present Gaze3P – the first large-scale dataset for studying user-perceived privacy using human eye gaze [^1] Gaze3P is available at https://collaborative-ai.org/research/datasets/Gaze3P/.
@@ 2. Preliminaries
where $\varepsilon\geq 0$ is the privacy budget, controlling the privacy loss. A small $\varepsilon$ means that $D$ and $D^{\prime}$ are (almost) not distinguishable given a set of outputs $S$. From an adversarial perspective, an adversary $\mathcal{A}$ challenged to distinguish $D$ and $D^{\prime}$ given an output set $S$ will output the dataset which is more likely, e.g. $D$ if $\Pr(D|S)\geq\frac{1}{2}\geq Pr(D^{\prime}|S)$. In the most extreme case of Eq. 1 we have $\Pr(M(D)\in S)=e^{\varepsilon}\Pr(M(D^{\prime})\in S)$ and hence $\Pr(D|S)=e^{\varepsilon}\Pr(D^{\prime}|S)=e^{\varepsilon}(1-\Pr(D|S))\Rightarrow\Pr(D|S)=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. Thus, the (absolute) advantage of an adversary is bounded by $\operatorname{adv}_{\mathcal{A}}\leq 2\frac{e^{\varepsilon}}{1+e^{\varepsilon}}-1=\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}$ (cf. Appendix C for more details).
@@ 3. Related Work
Eye Gaze and Privacy.  The results of the aforementioned works usually rely on user questionnaires. However, when using questionnaires, users often fail to follow their own privacy preferences (Orekondy et al., 2017; Plutzer, 2019; Mushure, 2017). In this paper, we propose to use human eye gaze (instead of questionnaires) to implicitly capture the dynamics of user-specific privacy perception. Eye-tracking data is already widely used to study human behaviour and cognition (Eckstein et al., 2017) and hypothesise that it can also be used as an indicator of user-perceived privacy. The idea to use gaze to detect privacy-sensitive situations is not completly new and has been explored in (Steil et al., 2019) where the users’ eye movements and first-person video were recorded using an egocentric (head-mounted) camera. However, (Steil et al., 2019) focuses on detecting privacy-sensitive situations rather than quantifying privacy perception. It also only features a small set of 17 participants in a free-viewing task and relies on recording and processing the scene imagery, which might break privacy (Orekondy et al., 2017). In this paper, we instead focus on the implicit and dynamic privacy perception feedback solely through gaze, without using the visual stimulus.
@@ 4. Gaze3P Dataset
[…]
@@
Stimuli.  We used a subset of the VISPR dataset (Orekondy et al., 2017) as our stimuli. Our dataset includes 1,000 images with corresponding private and safe (i.e. non-private) attributes. See Fig. 2 for sample images. The dataset contains 68 attributes categorised into nine attribute groups. The attributes were compiled according to the guidelines for the EU Data Protection Directive (GDPR) 95/46/EC, the US Privacy Act of 1974, and the data sharing rules in online social networks. The attribute categories are:
  • Personal information: e.g. age, gender, race, fingerprint, signature.
  • Documents: e.g. credit card, passport, national ID, driver’s license.
@@
  • Online activity: e.g. date/time of activity, username, password, email, online conversations.
@@
Participants.  We initially recruited $103$ participants through university mailing lists and notice boards. We had to exclude three participants due to calibration failures (Krafka et al., 2016). This resulted in a final group of $100$ participants ($37$ females and $63$ males). Participants were aged between $18$ and $35$ years, had different nationalities ($25$) and different (self-reported) privacy knowledge ($12$ experts). All participants had normal or corrected-to-normal vision. We refer to Appendix A for more demographic details.
@@
Compliance with the privacy and ethics guidelines.  The data was collected and processed according to the standards and guidelines of the ethical committee of the authors’ institution, with participants’ consent, remuneration, and pseudo-anonymisation procedures. In particular, the privacy and ethical guidelines of the Menlo Report (Dittrich and Kenneally, 2012) were satisfied.
@@ 5. Gaze3P Tasks
Our Gaze3P dataset enables new analyses that shed light on how users cognitively and behaviourally respond to privacy-relevant stimuli. More specifically, in this section we explore how well our gaze-based dataset Gaze3P is suited for two groups of learning tasks: (i) stimuli-based perceived privacy (SPP) tasks that infer how private a stimulus (e.g. image) is in Section 5.1 and (ii) user-based perceived privacy (UPP) tasks that infer informtion about the user (e.g. privacy expertise or identity) in Section 5.2. Each task entails learning a distinct mapping or pattern within the gaze data, such as predicting privacy ratings from gaze, classifying stimuli as private or non-private, or identifying user-specific privacy preferences based on demographic or behavioural features. We remark that we train all our models on annotated gaze data (i.e. with user-specified privacy ratings), such that they generalise to unseen users or stimuli. This not only facilitates the development of scalable and user-adaptive privacy-aware systems, but also provides empirical insights into the underlying mechanisms of perceived privacy.
In the following, we describe the two training tasks in detail and evaluate the performance of our trained models w.r.t standard baselines, namely decision tree (DT), support vector machine (SVM), logistic regression (LR), random forest (RF), K-nearest neighbour (KNN), convolutional neural networks (CNN), recurrent neural networks (RNN), and transformer (TF) models (cf. Table 1). To study individual stimuli, we focused on the free-viewing data of our dataset.
@@ 5.1. Stimuli-based Perceived Privacy (SPP)
SPP applications.  Before we discuss our results on tasks (1)–(4), we want to outline how stimuli-based perceived privacy results can be used in applications. For example, the quantified values in Tasks 1-3 (classified per attribute in Task 4, if needed) can be mapped to the corresponding parameters in privacy-preserving protocols. Such parameters can include $\varepsilon$ values in DP (Dwork et al., 2014; Dwork, 2006), model update perturbation or gradient clipping thresholds in federated learning (McMahan et al., 2017; Kairouz et al., 2021), and similar hand-picked parameters in K-anonymity (Samarati and Sweeney, 1998), L-diversity (Machanavajjhala et al., 2007), T-closeness (Li et al., 2006), privacy auctions (Ghosh and Roth, 2011; Zhang et al., 2020), synthetic data generation (Liu et al., 2022), etc. Other stimuli-specific insights are important for multiple applications, such as access control models (to define who can access which piece of information, e.g. attribute-based access control (ABAC) (Hu et al., 2015) which uses attributes for dynamic access control and human-in-the-loop privacy controls (e.g. Instagram’s ’Restrict’ feature (Parmelee and Roman, 2020) for controlling interactions, e-mail spam filters that allow manual corrections (Androutsopoulos et al., 2004), and marketing campaigns that respect the user’s perceived privacy preferences (Martin and Murphy, 2017)). We refer to Section 6 for sample applications.
SPP Baselines and Evaluation.  As baselines, we ran the aforementioned basic models. We evaluate the accuracy of the models with cross-validation and test sets to ensure that the model generalises well to unseen data. We also perform statistical hypothesis testing on the extracted eye-tracking features to examine whether the observed differences across experimental conditions or participant groups are statistically significant. These tests assess whether variations in metrics such as fixation duration, saccade amplitude, or pupil dilation are likely to reflect systematic effects rather than random noise or individual variability. To quantify the strength of evidence against the null hypothesis (i.e., that there is no meaningful difference between groups or conditions), we compute $p$-values. These values represent the probability of observing the given data, or something more extreme, under the assumption that the null hypothesis is true. A $p$-value below a conventional threshold (typically 0.05) is considered indicative of a statistically significant effect. Hence, we made the following key observations in relation to SPP Tasks 1-4:
@@
As shown in Table 1, our results demonstrate that gaze data can quantify perceived privacy since results exceed chance levels (0.5 for Task 1 and 0.14 for Tasks 2 and 3). We further analyse this in both person-specific and person-independent settings. In the person-specific setting, models are tested on individual participants, accounting for personal gaze behaviours, whereas in the person-independent approach, a generalizable model is trained across all participants and tested on 20% of each participant’s data. Results in Table 1 and Fig. 3 indicate significant variations in inter-user differences (person-independent, leading to lower accuracy), whereas intra-user features exhibit greater consistency (person-specific, leading to higher accuracy). This suggests that each user possesses a distinct and individualised perception of privacy.
@@
[…]
@@
For private attribute recognition, we compared our findings with the VISPR (Orekondy et al., 2017) evaluation. While they only input the stimuli, we replicated their models and metrics on our dataset and inputted the gaze data alone vs the gaze and stimuli. As shown in Table 2, when used in conjunction with stimuli, gaze data enhances the informational richness available to the model by providing additional features that complement the stimuli. However, even when gaze data is used in isolation, it remains beneficial while reducing the amount of shared data. Note that the models used are relatively basic for the sake of comparison with (Orekondy et al., 2017), and more advanced models could potentially enhance the results. Nonetheless, we observed that, as shown in Fig. 6 (left to right): (i) when using gaze as the only input, the models are able to identify objects that occupy a significant proportion of the image space while failing at identifying other subordinate attributes such as age or gender; (ii) when augmenting the stimulus with the gaze data, the models are able to identify more fine-grained details such as wedding rings and tattoos that the stimuli-alone version fails at (as long as the participants paid attention to such details), (iii) in all versions, models fail at identifying the relationship-based attributes (as also reported by (Orekondy et al., 2017)) since they require some reasoning capabilities that our basic models do not achieve.
Table 2. Evaluation of the private attribute recognition task as class-based mean average precision (C-MAP) (the average of the per-attribute average of the area under the Precision-Recall curve on all attributes), following (Orekondy et al., 2017)

### v2

@@ Abstract.
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXXccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
In this work, we explore a novel approach: The use of human eye gaze as an implicit and dynamic source of information on user-perceived privacy. In particular, our approach does not require a user to explicitly provide input or make active decisions, e,g, clicking a button or selecting a privacy setting. Naturally, the feedback derived from gaze is responsive to changing contexts, such as different content, tasks, or user states, i.e. as the user interacts with new stimuli, their gaze behaviour adapts, and our approach therefore continuously updates itself accordingly. Prior research has shown that eye gaze contains rich information about the user, such as identifiers (Cantoni et al., 2015), quasi-identifiers (Sammaknejad et al., 2017), confidential attributes (e.g. user activities (Steil and Bulling, 2015), attentive (Faber et al., 2018; Vertegaal et al., 2003) and cognitive states (Huang et al., 2016; Bulling and Roggen, 2011), or information about private situations (Steil et al., 2019)). Consequently, we try to answer our main research question RQ1:
[…]
@@
To this end, we present Gaze3P  (Gaze-based Prediction of user-Perceived Privacy)  – the first large-scale dataset for studying user-perceived privacy from the perspective of human eye gaze. Gaze3P includes gaze data (i.e. where, when, and how a person looks) of $100$ participants viewing $1000$ natural images showing different objects, some with private attributes (e.g. credit cards and medical history) as shown in Fig. 2. The dataset also provides user ratings of perceived privacy on a scale from 1 (very private) to 7 (very safe) for each image. The full dataset, including all annotations, will be made publicly available upon acceptance (cf. (Author, s) for the implementation). Using our new dataset, we then explore:
[…]
@@
We therefore present different tasks, focusing on the automatic prediction of users’ perceived privacy solely from gaze behaviour. These tasks correspond to privacy-related problems or objectives that an algorithm is trained to address using our collected data. Once trained, the algorithm can be used to generalise its learned solution to apply to the same class of problems for previously unseen individuals. Since it is not feasible to a priori determine which features can be reliably extracted from gaze data—nor whether potentially confounding factors can be effectively disentangled—we employed machine learning algorithms to automatically identify privacy-related patterns associated with each task. These tasks include: Stimuli-based Perceived Privacy (SPP) tasks to infer how private a stimulus (e.g. image) is and User-based Perceived Privacy (UPP) tasks to infer information about the user (e.g. privacy expertise or identity). Our ML models demonstrate that human eye gaze provides accurate predictions of perceived privacy.
Hence, we explore a third research question:
[…]
@@
We use the Gaze3P predictions of user-perceived privacy to optimise the parameters of differentially private mechanisms. Differentially private mechanisms obfuscate sensitive data samples such that only a limited amount of information about the private data can still be deduced from the obfuscated output of the mechanism. The exact amount of acceptable leakage depends on a privacy budget parameter $\varepsilon>0$, which determines the obfuscating noise added by the mechanism. If $\varepsilon$ is small, the privacy guarantee becomes stronger, but usually, the output of the mechanism is less accurate, and usability decreases. It is therefore important not to choose the privacy budget $\varepsilon$ too small, i.e., to only add the minimal amount of noise that guarantees a target privacy level. The optimisation of DP-parameters has therefore seen much attention in recent years (Jorgensen et al., 2015; Boenisch et al., 2022; Niu et al., 2021).
@@
Our new dataset Gaze3P and the resulting ML model predictions provide a new way to determine $\varepsilon$, which reflects a user’s perceived privacy. Depending on the actual use case, we propose different mappings from perceived privacy levels to $\varepsilon$-values. We evaluate how each mapping affects the utility of the obfuscated output dataset and show that our gaze-based approach outperforms previous work.
@@
    We present Gaze3P – the first large-scale dataset for studying user-perceived privacy using human eye gaze, providing a dynamic and implicit user feedback.
@@ 2. Preliminaries
By transforming the continuous, high-dimensional gaze signals into quantitative features—such as fixations, saccades, and pupil diameters—machine learning (ML) models are provided with informative inputs that capture essential characteristics of user interaction or cognitive states. These features are then passed into ML algorithms, enabling the models to learn underlying statistical patterns or associations within the data. This learned structure allows the models to perform specific tasks such as classification, regression, or clustering. Ultimately, this feature-to-model pipeline allows ML systems to generalize from training data and make accurate, data-driven inferences about new, unseen inputs.
@@
where $\varepsilon\geq 0$ is the privacy budget, controlling the privacy loss. A small $\varepsilon$ means that $D$ and $D^{\prime}$ are (almost) not distinguishable given a set of outputs $S$. From an adversarial perspective, an adversary $\mathcal{A}$ challenged to distinguish $D$ and $D^{\prime}$ given an output set $S$ will output the dataset which is more likely, e.g. $D$ if $\Pr(D|M(D)\in S)\geq\frac{1}{2}\geq\Pr(D|M(D^{\prime})\in S)$. In the most extreme case of Eq. 1 we have $\Pr(M(D)\in S)=e^{\varepsilon}\Pr(M(D^{\prime})\in S)$ and hence $\Pr(D|M(D)\in S)=e^{\varepsilon}\Pr(D^{\prime}|M(D)\in S)=e^{\varepsilon}(1-\Pr(D|M(D)\in S))\Rightarrow\Pr(D|M(D)\in S)=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. Thus, the (absolute) advantage of an adversary is bounded by $\operatorname{adv}_{\mathcal{A}}\leq 2\frac{e^{\varepsilon}}{1+e^{\varepsilon}}-1=\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}$ (cf. Appendix E for more details).
@@ 3. Related Work
Eye Gaze and Privacy Perception.  The results of the aforementioned works usually rely on user questionnaires. However, when using questionnaires, users often fail to follow their own privacy preferences (Orekondy et al., 2017; Plutzer, 2019; Mushure, 2017). In this paper, we propose to use human eye gaze (instead of questionnaires) to implicitly capture the dynamics of user-specific privacy perception. Eye-tracking data is already widely used to study human behaviour and cognition (Eckstein et al., 2017; Carter and Luke, 2020; Hessels et al., 2025; Hooge et al., 2025; Nyström et al., 2025; Niehorster et al., 2025). Prior research has demonstrated that gaze patterns can reflect cognitive processes such as risk perception and habituation (Anderson et al., 2016; Guerra et al., 2023a, 2024). For instance, studies have shown that increased risk is often associated with longer fixations and heightened visual scanning, indicating deeper cognitive engagement (Toma et al., 2023; Mormann and Frydman, 2016; Harrison and Swarthout, 2019). Conversely, habituation to repeated stimuli can lead to reduced gaze variability and decreased attention, even in the presence of sensitive information—a challenge for maintaining consistent privacy awareness (Anderson et al., 2016). Gaze3P builds on these insights and hypothesises that gaze can also be used as an indicator of user-perceived privacy. The idea to use gaze to detect privacy-sensitive situations is not completely new and has been explored in (Steil et al., 2019), where the users’ eye movements and first-person video were recorded using an egocentric (head-mounted) camera. However, (Steil et al., 2019) focuses on detecting privacy-sensitive situations rather than quantifying privacy perception. It also only features a small set of 17 participants in a free-viewing task and relies on recording and processing the scene imagery, which might break privacy (Orekondy et al., 2017). In this paper, we instead focus on the implicit and dynamic privacy perception feedback solely through gaze.
Other Privacy-Related Gaze Applications.  Prior works at the intersection of eye tracking and privacy mainly focus on (i) eye-based authentication (Luo et al., 2020; Liebers and Schneegass, 2020), (ii) privacy considerations and guidelines (Gressel et al., 2023; Katsini et al., 2020), (iii) secure AR/VR applications (Luo et al., 2020; Li et al., 2021; Bozkir et al., 2023; David-John et al., 2021a, b, 2022), (iv) UI design for secure interactions (Katsini et al., 2020), (v) secure gaze data sharing (Elfares et al., 2022, 2023, 2025; Bozkir et al., 2021; David-John et al., 2021a, b, 2022), and (vii) information leakage and attacks on gaze data (Wang et al., 2024; Chen et al., 2018; Elfares et al., 2023; Sönnichsen et al., 2025). Apart from the different focus, Gaze3P, also differs in (i) the purpose of gaze, i.e., some prior work (Wang et al., 2024; Luo et al., 2020; Li et al., 2021) uses gaze for explicit interaction (e.g., authentication (Luo et al., 2020) or gaze rays in VR (Li et al., 2021)) while Gaze3P  uses gaze as an implicit (passive) signal to infer perceived privacy, (ii) unlike prior work, we focus on the cognitive aspect of gaze in privacy, beyond the gaze location estimation (Wang et al., 2024; Chen et al., 2018), (iii) privacy positioning, i.e. Gaze3P is user-centric and proactive (helping users protect their own privacy through implicit gaze behaviour) while prior works either focus on system-enforced protection for eye tracking (Elfares et al., 2022, 2023, 2025; Li et al., 2021; David-John et al., 2021a, b, 2022), user authentication (Luo et al., 2020; Liebers and Schneegass, 2020), or demonstrating privacy vulnerabilities (Katsini et al., 2020; Sönnichsen et al., 2025).
@@ 4. Gaze3P Dataset
Eye tracker.  For gaze data collection, we used an Eyelink eye tracker that provides binocular gaze data at a sampling rate of $2\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. As is common practice in laboratory eye tracking studies, we used a chin rest to stabilise participants’ heads. Images were shown on a computer screen with a resolution of 1920x1080 pixels and a size of $545\text{\,}\mathrm{m}\mathrm{m}$ x $303\text{\,}\mathrm{m}\mathrm{m}$. The eye-to-screen distance was $700\text{\,}\mathrm{m}\mathrm{m}$. The proportion of the calibrated area was set to 0.63 x 0.88 to stay within the trackable range of the system with an HV13-13 (horizontal/vertical 13 targets) calibration type for better spatial accuracy across the entire screen. The recorded gaze data was then processed into fixations, saccades, and pupil information (cf. Section 2). More details and visualisations can be found in Appendix A.
@@
Stimuli.  We randomly sampled a subset of images from the VISPR dataset (Orekondy et al., 2017) as our stimuli since it is the only publicly-available dataset that contains privacy-related attributes. See Fig. 2 for sample images. The dataset contains 68 attributes categorised into nine attribute groups. The attributes were compiled according to the guidelines for the EU Data Protection Directive (GDPR) 95/46/EC, the US Privacy Act of 1974, and the data sharing rules in online social networks. It further includes reliable and consistent attribute annotations by letting multiple annotators follow detailed labelling guidelines. We ensured a balanced distribution across the annotated privacy attributes. This stratified sampling approach was employed to maximise attribute coverage and mitigate sampling bias, given the practical constraint on the number of stimuli each participant could reasonably view during the experiment. Despite this limitation, the resulting dataset remains relatively large and representative, supporting robust analysis of visual privacy perception across a diverse range of attributes. Our dataset includes 1,000 images with corresponding private and safe (i.e. non-private) attributes. The attribute categories in VISPR (Orekondy et al., 2017) are:
  • Personal information: e.g. age, gender, fingerprint, signature.
  • Documents: e.g. credit card, passport, national ID.
@@
  • Online activity: e.g. date/time of activity, username, password.
@@
Participants.  We initially recruited $103$ participants through university mailing lists and notice boards. We had to exclude three participants due to calibration failures (Krafka et al., 2016). This resulted in a final group of $100$ participants ($37$ females and $63$ males). Participants were aged between $18$ and $35$ years, had different nationalities ($25$) and different (self-reported) privacy knowledge ($12$ experts) [^1] All experts are MSc or PhD holders of information security degrees.. All participants had normal or corrected-to-normal vision. We refer to Appendix A for more demographic details.
@@
Compliance with the privacy and ethics guidelines.  The data was collected and processed according to the standards, guidelines, and approval of the ethical committee of the authors’ institution, with participants’ consent, remuneration, and pseudo-anonymisation procedures. In particular, the privacy and ethical guidelines of the Menlo Report (Dittrich and Kenneally, 2012) were satisfied.
@@ 5. Gaze3P Tasks
Our Gaze3P dataset enables new analyses that shed light on how users cognitively and behaviourally respond to privacy-relevant stimuli. Given the difficulty in determining beforehand which features can be meaningfully extracted from gaze data and whether confounding variables can be effectively separated, we adopted a machine learning-based approach to automatically uncover privacy-related patterns corresponding to each task. This method enables us to objectively evaluate our hypothesis: that gaze behaviour may function as a proxy for perceived privacy.
More specifically, in this section, we explore how well our gaze-based dataset Gaze3P is suited for two groups of learning tasks: (i) stimuli-based perceived privacy (SPP) tasks that infer how private a stimulus (e.g. image) is in Section 5.1 and (ii) user-based perceived privacy (UPP) tasks that infer information about the user (e.g. privacy expertise or identity) in Section 5.2. Each task entails learning a distinct mapping or pattern within the gaze data, such as predicting privacy ratings from gaze, classifying stimuli as private or non-private, or identifying user-specific privacy preferences based on demographic or behavioural features. For implementation, we used a basic Scikit-learn framework with default parameter settings. This provided a standardised and unbiased baseline for model training, avoiding manual feature selection or tuning that could skew results or introduce overfitting. It is important to note that all models in this study are trained on annotated gaze data, where user-specified privacy ratings serve as ground-truth labels to facilitate supervised learning and enable robust evaluation of model performance. The purpose of this training procedure is to allow the model to learn association patterns for each task. During testing and deployment, however, only the raw gaze data is provided as input to the model—without any accompanying user-specified ratings. This setup reflects a practical application scenario, wherein the model is expected to generalise to previously unseen users or stimuli and autonomously infer privacy-related judgments based solely on gaze behaviour. This not only facilitates the development of scalable and user-adaptive privacy-aware systems but also provides empirical insights into the underlying mechanisms of perceived privacy.
In the following, we describe the two training tasks in detail and evaluate the performance of our trained models w.r.t standard baselines, namely decision tree (DT), support vector machine (SVM), logistic regression (LR), random forest (RF), K-nearest neighbour (KNN), and transformer (TF) models (cf. Table 1). To study individual stimuli, we focused on the free-viewing data of our dataset.
@@ 5.1. Stimuli-based Perceived Privacy (SPP)
SPP applications.  Before we discuss our results on tasks (1)–(4), we want to outline how stimuli-based perceived privacy results can be used in applications. For example, the quantified values in Tasks 1-3 (classified per attribute in Task 4, if needed) can be mapped to the corresponding parameters in privacy-preserving protocols. Such parameters can include $\varepsilon$ values in DP (Dwork et al., 2014; Dwork, 2006), model update perturbation or gradient clipping thresholds in federated learning (McMahan et al., 2017; Kairouz et al., 2021), and similar hand-picked parameters in K-anonymity (Samarati and Sweeney, 1998), L-diversity (Machanavajjhala et al., 2007), T-closeness (Li et al., 2006), privacy auctions (Ghosh and Roth, 2011; Zhang et al., 2020), synthetic data generation (Liu et al., 2022a), etc. Other stimuli-specific insights are important for multiple applications, such as access control models (to define who can access which piece of information, e.g. attribute-based access control (ABAC) (Hu et al., 2015) which uses attributes for dynamic access control and human-in-the-loop privacy controls (e.g. Instagram’s ’Restrict’ feature (Parmelee and Roman, 2020) for controlling interactions, e-mail spam filters that allow manual corrections (Androutsopoulos et al., 2004), and marketing campaigns that respect the user’s perceived privacy preferences (Martin and Murphy, 2017)). We refer to Section 6 for sample applications and to Appendix C for further details.
SPP Baselines and Evaluation.  As baselines, we ran the aforementioned basic models. We evaluate the accuracy of the models with cross-validation and test sets to ensure that the model generalises well to unseen data. Implementation details can be found in Appendix B. We also perform statistical hypothesis testing on the extracted eye-tracking features to examine whether the observed differences across experimental conditions or participant groups are statistically significant. These tests assess whether variations in metrics such as fixation duration, saccade amplitude, or pupil dilation are likely to reflect systematic effects rather than random noise or individual variability. To quantify the strength of evidence against the null hypothesis (i.e., that there is no meaningful difference between groups or conditions), we compute $p$-values. These values represent the probability of observing the given data, or something more extreme, under the assumption that the null hypothesis is true. A $p$-value below a conventional threshold (typically 0.05) is considered indicative of a statistically significant effect. Hence, we made the following key observations in relation to SPP Tasks 1-4:
@@
Gaze Saliency Maps
DT 0.45 0.17 0.21
SVM 0.52 0.30 0.31
LR 0.57 0.29 0.30
RF 0.52 0.23 0.27
KNN 0.50 0.18 0.18
TF 0.46 0.29 0.32
Residual Gaze Maps
DT 0.23 0.11 0.13
SVM 0.21 0.15 0.22
LR 0.19 0.17 0.08
RF 0.27 0.19 0.21
KNN 0.09 0.10 0.23
TF 0.31 0.18 0.19
@@
Table 1 shows the accuracy of the aforementioned models if all (meta)data is passed to the ML model. Our results demonstrate that gaze data can quantify perceived privacy since results exceed chance levels (0.5 for Task 1 and 0.14 for Tasks 2 and 3). We further analyse this in both person-specific and person-independent settings. In each case, models were trained on 80% of the data and tested on the remaining 20% with cross-validation. In the person-specific setting, models are tested on 20% of individual participants, accounting for personal gaze behaviours (with only one sample at a time independent of any user history), whereas in the person-independent approach, a generalizable model is trained across all participants and tested on 20% of each participant’s data. Results in Table 1 and Fig. 3 indicate significant variations in inter-user differences (person-independent, leading to lower accuracy), whereas intra-user features exhibit greater consistency (person-specific, leading to higher accuracy). This suggests that each user possesses a distinct and individualised perception of privacy.
@@
In general, fixation duration increases on private cues with fewer fixations on less private regions ($H_{0}$: ’There is no difference in the distribution of fixation durations across different stimuli regions’ tested with Kruskal-Wallis H-test and Dunn’s test with Bonferroni correction for multiple comparisons with the different privacy levels, with $p\text{-}value=$0.04$<$0.05$\$ and Epsilon-squared effect size for Kruskal-Wallis H-test of $0.05$). This can also be seen in Fig. 3. It further suggests that less fine-grained tracking is sufficient (i.e. relying on fixations alone). This was further supported by downsampling the gaze data to $30\text{\,}\mathrm{H}\mathrm{z}$ ($30\text{\,}\mathrm{F}\mathrm{P}\mathrm{S}$) to simulate commodity-level standard webcams, as shown in Table 1-downsampling. The results, again, suggest that high-resolution gaze data (Table 1-person-independent) may not be necessary for effective privacy perception prediction, facilitating a more widespread implementation and enhancing accessibility and usability for a broader audience  [^3] As this study represents exploratory research, we employed a high-resolution, constrained experimental setup. This choice was motivated by the initial uncertainty regarding whether a lower-resolution configuration would yield significant results and to what extent fine-grained details in the data would be necessary to capture relevant effects. The high-resolution setup ensured maximal data fidelity, allowing for comprehensive observation of potentially subtle phenomena during the early stages of research..
@@
For private attribute recognition, as shown in Table 2, we compared our findings with the VISPR (Orekondy et al., 2017) evaluation as well as other SOTA models. While (Orekondy et al., 2017) only input the stimuli, we replicated their models and metrics on our dataset and inputted the gaze data alone vs the gaze and stimuli. Nonetheless, we observed that, as shown in Fig. 6 (left to right): (i) when using gaze as the only input, the models are able to identify objects that occupy a significant proportion of the image space while failing at identifying other subordinate attributes such as age or gender; (ii) when augmenting the stimulus with the gaze data, the models are able to identify more fine-grained details such as wedding rings and tattoos that the stimuli-alone version fails at (as long as the participants paid attention to such details), (iii) in all versions, models fail at identifying the relationship-based attributes (as also reported by (Orekondy et al., 2017)) since they require some reasoning capabilities that our basic models do not achieve.
Table 2. Evaluation of the private attribute recognition task as class-based mean average precision (C-MAP) (the average of the per-attribute average of the area under the Precision-Recall curve on all attributes), following (Orekondy et al., 2017)’s models (CaffeNet, GoogleNet, and ResNet) and more recent SOTA models (Multimodal fusion model: GazeFormer (Mondal et al., 2023), Cross-modal transformer: ViLT (Kim et al., 2021), and self-supervised contrastive model: CLIP (Radford et al., 2021))
@@
SOTA GazeFormer 64.13 40.82 92.73
SOTA ViLT 66.18 41.59 89.47
SOTA CLIP 61.00 35.06 78.56